Find all solutions of the equation below on the interval [0,2pi). Enter your answer as a comma-separated list of values or enter none if there is no solution.
cos theta cos 3 theta - sin theta sin 3 theta = 0
Answer: theta =

Answer by stanbon(75887)

Find all solutions of the equation below on the interval [0,2pi). Enter your answer as a comma-separated list of values or enter none if there is no solution.
cos theta cos 3 theta - sin theta sin 3 theta = 0
------
Formula Form: cos(x+y) = cos(x)cos(y) - sin(y)sin(x)
----------
Answer: Your problem has that form
---
cos(3t+t) = 0
----
cos(4t) = 0
----
Equation:
4t = pi/2 or (3/2)pi
---------------------------
t = pi/8 or t = (3/8)pi
===========================
Cheers,
Stan H.
==================